SYSTEMA 3x +2y = 8 5x-3y = 7 GRAPHING METHOD SUBSTITUTION METHOD ELIMINATION METHOD Collaborative Learning Activities
Answer #1

3x+2y = 8 +0 5x-3y 27 @ Graphing method 147 @ It 5 / -5 1 5 substitution method from ①, ey=8-34 ... >yo - (832) putting the y